OpenStudy (anonymous):

The school cafeteria sells two kinds of wraps: vegetarian and chicken. The vegetarian wrap costs $1.00 and the chicken wrap costs $1.80. Today they made $98.80 from the 70 wraps sold. How many of the wraps sold were vegetarian?

OpenStudy (anonymous):

let x=vegetarian y=chicken equations: (1) x+y=70 (2) x+1.8y=98.8

OpenStudy (anonymous):

how many veggie wraps was sold

OpenStudy (anonymous):

....